OpenStudy (anonymous):

Simplify the expression 3(5y + 6) - 4

OpenStudy (anonymous):

Use the distributive property: x(y + z) = xy + xz for the first part

OpenStudy (anonymous):

Here is an example: 4(2y + 3) = 4*2y + 4*3 = 8y +12

OpenStudy (godlovesme):

 make sense @helpmeplz11